SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Feb.14, 2006)

Attendees: Rainer Bartoldus, Fang Fang, Selina Li, Rahmat Rahmat, Dan Walker

Operations

  • Investigation of the vacuum leak for PEP is still undergoing. BaBar has been taking data at lumi of about 4-5x10^33 for the past week or so. Trigger is in stable operation.
  • The DCH front-end FEX mode was enabled for run 62240 and subsequent runs.

EMT

  • On Sunday (Feb.12) at 17:50:31, there is an entry in the log book indicating that the EMT crate has 0x2 damage (dropped contributions). Rainer explained to us that sometimes we get dropped contribution because EMT didn't come on time to contribute to an event. After talking to Chris O'Grady (after this meeting), we found out that the EMT ROM seemed to be hung and it had happened at least once before. Chris suggested that it was a hardware problem with the ODF (or a software problem which could be due to memory leak, but it doesn't look very likely). He started to keep a record of this problem.
  • Both Selina and Dan seem to have problem accessing the EMT deadtime archive. It looked like the information was lost for some period of time. Dan will contact the appropriate expert(s) to fine out what happened.
  • The turn-on-curve (see example of the fastmon plots) for EMT seemed to work in Dan's own release, but not in the postproc for making the ps archive (plots after the efficiency for turn-on-curve were not shown). Dan is looking into this problem.

DCT/DCZ

  • The MPIC Spurious Interrupt and out-of-frame problem is discussed in HN posted. Rainer asked what the beam condition was like for the worst case on Jan.28. Looking back, it looked like we ran at very low lumi for most of that day, and there were some high background period as well.

DCT online

  • Fang linked in the DCZ online lib with debug statement put in to tell us which TSF element/board is out-of-frame in the ROM logfile (on Friday and again on Monday).
  • Rainer and Fang will get together this Friday to go over the DCZ online code with removal of the dependency of the test stand codes.

L3

  • While the bug that Rainer fixed last week had a rather dramatic effect on beam-gas events, it had a negligible effect (10^-5 level) on overall L1 Accepts (and physics filters). So it is safe to be installed without much adaptation. It's a good idea to implement since it will remove bias in L3 tracking.

Data Quality and Monitoring

  • The Trigger data from last week are okay, including the test runs with the new DCH FEX mode implemented.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.